How many liters of a 10% solution should be mixed with a 40% sol to obtain 50 liters of a 25% sol?
Answer by robertb(5830)

0.10x + 0.40(50-x) = 0.25*50,
0.10x + 20 - 0.40x = 12.5,
-0.30x = -7.5,
x = 25, # liters of 10% solution.